# InfinitePOD: Building Datacenter-Scale High-Bandwidth Domain for LLM with Optical Circuit Switching Transceivers

### v2

@@ 8. Related Work
HBD Architectures. HBDs are crucial for enabling communication intensive parallelism strategies (TP/EP) for LLM training. NVIDIA DGX SuperPOD (NVIDIA, 2024a) and GB200 NVL series (NVIDIA, 2024b) use any-to-any electrical switching, delivering high performance but suffering from high costs, scalability limitations, and fragmentation. In contrast, direct interconnect HBDs like Dojo (Tesla, 2024), TPUv3 (Jouppi et al., 2020), and SiP-Ring (Khani et al., 2021) improve scalability but have a large fault explosion radius. TPUv4 (Jouppi et al., 2023) and TPUv5p (Google, 2024) attempts a middle ground but still lacks full node-level fault isolation. InfinitePOD introduces a novel architecture that reduces cost, improves scalability, minimizes fragmentation, enhances fault isolation, and dynamically supports TP.
AI DCN Architectures. MegaScale (Jiang et al., 2024a) and Meta’s (Gangidi et al., 2024b) AI DC use Clos-based topologies, while Rail-Optimized (NVIDIA, 2021) and Rail-Only (Wang et al., 2024) architectures optimize for LLM traffic patterns. Alibaba HPN (Qian et al., 2024) enhances fault tolerance with a dual-plane design. InfinitePOD is compatible with all of them on LLM-training.
@@ 9. Conclusion
In this paper, we propose InfinitePOD, a novel HBD design that supports datacenter scale, dynamic TP group size and near-ideal fault explosion radius. InfinitePOD is built upon a novel design of optical transceivers interaged with SiPh-based OCSTrx, a reconfigurable K-Hop Ring topology and a HBD-DCN orchestration algorithm to leverage the capabilities of the new hardware. Using real fault trace of 3K GPU cluster and the in-house simulator, we demonstrate that InfinitePOD achieves GPU utilization close to the ideal model during faults, delivers superior cost and energy efficiency compared to existing designs, and provides effective control over cross-ToR DCN traffic. We believe InfinitePOD provides an efficient scaling solution for HBD, which offers new insights for the infrastructure of the next generation Trillion-Parameters LLM training.
@@ Appendix A Production Fault Trace
The production fault trace was collected from an 8-GPU node pretrain cluster with 2880 GPUs over a period of 160 days. The trace includes details such as fault start time, fault end time, and the ID of the faulty node. Figure 14(a) and Figure 14(b) provide a macro-level overview of the production fault trace. On average, the ratio of faulty 8-GPU nodes at any given time is $3.83\%$, with a p99 value of $7.22\%$.
@@
$P_{fault}(8\text{-GPU})=1-(1-p)^{8}=3.83\%.$
From this, we derive $p=0.49\%$. The fault rate for a 4-GPU node is then:
$P_{fault}(4\text{-GPU})=1-(1-p)^{4}=1.93\%.$
@@
$=\frac{1\times 1.93\%}{3.83\%}=50.39\%$
Thus, whenever a fault occurs in an 8-GPU node in the original trace, each of the two corresponding 4-GPU nodes at the same location has a $50.39\%$ probability of fault. This method is used to convert the traces.
@@ Appendix B GPT-MoE Architecture
Appendix C Theoretical analysis of wasted GPU ratio for InfinitePOD
The count of backup lines as $22K-$ 2 will significantly influence the fault tolerance of InfinitePOD. We use the expectation of waste ratio caused by GPU failure and fragmentation problem to evaluate this design, the result is shown in Table 7.
For one single working server in the middle of line, the count of breakpoints $B$ on its two sides has the expectation as:
@@ Appendix C Theoretical analysis of wasted GPU ratio for InfinitePOD
Where $P_{s}$ is the fail probability of GPU server, and $\eta$ is count of servers. The expectation of breakpoints count is:
Once the distance between one server and the tail of line is $\alpha<K$, it will connect to all servers between itself and the last one, so there will be no breakpoints on this side, and the expectation of breakpoints count is less than servers in the middle of line. Then, for any server in the line topology:
@@
When the distance between two servers is $\beta\geq K$, the breakpoints among them can be calculated as independent. Once the distance $\beta<K$, as all servers in this range are connected to these two servers, there will be no breakpoints between them. So, the expectation is less than two independent servers. Then,
@@
For a LLM job which require a ring communication size (TP .etc) as $N_{t}$, InfinitePOD will cut the whole line topology into several sub lines with the length of $N_{t}/R$. Once InfinitePOD is cutting a new sub line from the remaining servers in the line, all $N_{t}$ GPU will be wasted when one break point exist in the middle of this sub line required, shown in Figure 15. Then the expectation for waste GPU caused by one single break point is:
@@
Figure 15. Break point can cause server waste compare to ideal situation.
@@
So the expectation of wasted GPU for a servers cluster with $N_{s}$ GPU servers is:
@@
In our trace for a 160 days long pre-train job on 10K-GPU, the p99 failure rate for 8-card machines is 7%. If a TP32 jobs is running on InfinitePOD, we can get the upper bond for waste ratio expectation for various configuration in Table 7.
@@
R=4 $7.35\%$ $0.26\%$ $9.00\times 10^{-4}$
R=8 $27.4\%$ $1.92\%$ $0.13\%$
Table 7. Upper bond for waste ratio expectation of GPU, where GPU failure rate is 0.875% and X is 32
As shown in the table, for 4 GPU server ($R=4$) 3 bundles ($K=3$) design, the additional waste of GPU is less than 0.26%, while the waste ratio for $R=8,K=4$ is less than 0.13%. This is sufficient for production clusters.
@@ Appendix D Orchestration For Fat-Tree
  • $S$: ordered subset, represents nodes, $\forall u\in S,u\in S_{all}$. Adjacent elements in $S$ are also adjacent from the perspective of the InfinitePOD topology.
@@
  • $InfHBD=<S,E>$: the topology of InfinitePOD as an undirected graph.
@@
  • $G_{deploy}=<S_{deploy},E_{deploy}>$: deployed topology. After applying the deployment strategy, the topology from the perspective of InfinitePOD is described as follows: $S_{\text{deploy}}$ is an ordered set where adjacent elements correspond to adjacent nodes in InfinitePOD, and $E_{\text{deploy}}$ represents the connections between nodes.
@@ Appendix F Detailed Cost and power consumption Analysis
In this section, Table 8 provides a detailed description of the quantity, cost, bandwidth, and power consumption of the interconnect components in various network architectures, including Google TPUv4 (Jouppi et al., 2023), NVIDIA GB200 NVL series (NVIDIA, 2024b), Alibaba HPN(Qian et al., 2024), and InfinitePOD.
@@
InfinitePOD($K=2$) with 4 GPU, bandwidth 800GBps/GPU
@@
InfinitePOD($K=3$) with 4 GPU, bandwidth 800GBps/GPU

### v3

@@ 8. Related Work
HBD Architectures. HBDs are crucial for enabling communication intensive parallelism strategies (TP/EP) for LLM training. NVIDIA DGX SuperPOD (NVIDIA, 2024a) and GB200 NVL series (NVIDIA, 2024b) use any-to-any electrical switching, delivering high performance but suffering from high costs, scalability limitations, and fragmentation. In contrast, direct interconnect HBDs like Dojo (Tesla, 2024), TPUv3 (Jouppi et al., 2020), and SiP-Ring (Khani et al., 2021) improve scalability but have a large fault explosion radius. TPUv4 (Jouppi et al., 2023) and TPUv5p (Google, 2024) attempts a middle ground but still lacks full node-level fault isolation. InfiniteHBD introduces a novel architecture that reduces cost, improves scalability, minimizes fragmentation, enhances fault isolation, and dynamically supports TP.
AI DCN Architectures. MegaScale (Jiang et al., 2024a) and Meta’s (Gangidi et al., 2024b) AI DC use Clos-based topologies, while Rail-Optimized (NVIDIA, 2021) and Rail-Only (Wang et al., 2024) architectures optimize for LLM traffic patterns. Alibaba HPN (Qian et al., 2024) enhances fault tolerance with a dual-plane design. InfiniteHBD is compatible with all of them on LLM-training.
@@ 9. Conclusion
In this paper, we propose InfiniteHBD, a novel HBD design that supports datacenter scale, dynamic TP group size and near-ideal fault explosion radius. InfiniteHBD is built upon a novel design of optical transceivers interaged with SiPh-based OCSTrx, a reconfigurable K-Hop Ring topology and a HBD-DCN orchestration algorithm to leverage the capabilities of the new hardware. Using real fault trace of 10K GPU cluster and the in-house simulator, we demonstrate that InfiniteHBD achieves GPU utilization close to the ideal model during faults, delivers superior cost and energy efficiency compared to existing designs, and provides effective control over cross-ToR DCN traffic. We believe InfiniteHBD provides an efficient scaling solution for HBD, which offers new insights for the infrastructure of the next generation Trillion-Parameters LLM training.
@@ Appendix A Production Fault Trace
The production fault trace was collected from an 8-GPU node pretrain cluster with 10K GPUs over a period of 348 days. The trace includes details such as fault start time, fault end time, and the ID of the faulty node. Figure 14(a) and Figure 14(b) provide a macro-level overview of the production fault trace. On average, the ratio of faulty 8-GPU nodes at any given time is $2.66\%$, with a p99 value of $7.92\%$.
@@
$P_{fault}(8\text{-GPU})=1-(1-p)^{8}=2.66\%.$
From this, we derive $p=0.34\%$. The fault rate for a 4-GPU node is then:
$P_{fault}(4\text{-GPU})=1-(1-p)^{4}=1.34\%.$
@@
$=\frac{1\times 1.34\%}{2.66\%}=50.38\%$
Thus, whenever a fault occurs in an 8-GPU node in the original trace, each of the two corresponding 4-GPU nodes at the same location has a $50.38\%$ probability of fault. This method is used to convert the traces.
@@ Appendix B GPT-MoE Architecture
Appendix C Theoretical analysis of wasted GPU ratio for InfiniteHBD
The count of backup lines as $22K-$ 2 will significantly influence the fault tolerance of InfiniteHBD. We use the expectation of waste ratio caused by GPU failure and fragmentation problem to evaluate this design, the result is shown in Table 7.
For one single working node in the middle of line, the count of breakpoints $B$ on its two sides has the expectation as:
@@ Appendix C Theoretical analysis of wasted GPU ratio for InfiniteHBD
Where $P_{s}$ is the fail probability of GPU node, and $\eta$ is count of nodes. The expectation of breakpoints count is:
Once the distance between one node and the tail of line is $\alpha<K$, it will connect to all nodes between itself and the last one, so there will be no breakpoints on this side, and the expectation of breakpoints count is less than nodes in the middle of line. Then, for any node in the line topology:
@@
When the distance between two nodes is $\beta\geq K$, the breakpoints among them can be calculated as independent. Once the distance $\beta<K$, as all nodes in this range are connected to these two nodes, there will be no breakpoints between them. So, the expectation is less than two independent nodes. Then,
@@
For a LLM job which require a ring communication size (TP .etc) as $N_{t}$, InfiniteHBD will cut the whole line topology into several sub lines with the length of $N_{t}/R$. Once InfiniteHBD is cutting a new sub line from the remaining nodes in the line, all $N_{t}$ GPU will be wasted when one break point exist in the middle of this sub line required, shown in Figure 15. Then the expectation for waste GPU caused by one single break point is:
@@
Figure 15. Break point can cause node waste compare to ideal situation.
@@
So the expectation of wasted GPU for a nodes cluster with $N_{s}$ GPU nodes is:
@@
In our trace for a 348 days long pre-train job on 10K-GPU, the p99 failure rate for 8-GPU nodes is 7.92%. If a TP32 jobs is running on InfiniteHBD, we can get the upper bond for waste ratio expectation for various configuration in Table 7.
@@
R=4 $9.22\%$ $0.37\%$ $1.52\times 10^{-4}$
R=8 $30.11\%$ $2.38\%$ $0.19\%$
Table 7. Upper bond for waste ratio expectation of GPU, where GPU failure rate is 1.03% and $N_{t}$ is 32
As shown in the table, for 4-GPU node ($R=4$) 3 bundles ($K=3$) design, the additional waste of GPU is less than 0.37%, while the waste ratio for $R=8,K=4$ is less than 0.19%. This is sufficient for production clusters.
@@ Appendix D Orchestration For Fat-Tree
  • $S$: ordered subset, represents nodes, $\forall u\in S,u\in S_{all}$. Adjacent elements in $S$ are also adjacent from the perspective of the InfiniteHBD topology.
@@
  • $InfHBD=<S,E>$: the topology of InfiniteHBD as an undirected graph.
@@
  • $G_{deploy}=<S_{deploy},E_{deploy}>$: deployed topology. After applying the deployment strategy, the topology from the perspective of InfiniteHBD is described as follows: $S_{\text{deploy}}$ is an ordered set where adjacent elements correspond to adjacent nodes in InfiniteHBD, and $E_{\text{deploy}}$ represents the connections between nodes.
@@ Appendix F Detailed Cost and power consumption Analysis
In this section, Table 8 provides a detailed description of the quantity, cost, bandwidth, and power consumption of the interconnect components in various network architectures, including Google TPUv4 (Jouppi et al., 2023), NVIDIA GB200 NVL series (NVIDIA, 2024b), Alibaba HPN(Qian et al., 2024), and InfiniteHBD.
@@
InfiniteHBD($K=2$) with 4 GPU, bandwidth 800GBps/GPU
@@
InfiniteHBD($K=3$) with 4 GPU, bandwidth 800GBps/GPU